# Loop Feynman integration on a quantum computer

### v1

@@ Acknowledgements.
We acknowledge Matteo Robbiati for providing help and support in executing the QNN on the Qibo hardware and Stefano Carrazza for providing us access to the quantum device. We acknowledge financial support from the Spanish Government (Agencia Estatal de Investigación MCIN/AEI/ 10.13039/501100011033) Grant No. PID2020-114473GB-I00, and Generalitat Valenciana Grants No. PROMETEO/2021/071 and ASFAE/2022/009 (Planes Complementarios de I+D+i, NextGenerationEU). This work has also been funded by the Ministry of Economic Affairs and Digital Transformation of the Spanish Government through the QUANTUM ENIA - QUANTUM SPAIN project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda, and by CSIC Interdisciplinary Thematic Platform (PTI+) on Quantum Technologies (PTI-QTEP+). JML acknowledges financial support from Generalitat Valenciana (Grant No. ACIF/2021/219) and CSIC (Grant No. IMOVE23124). LC is supported by Generalitat Valenciana GenT Excellence Programme (CIDEGENT/2020/011) and ILINK22045. MG and SV are supported by CERN through the CERN QTI. Access to the IBM Quantum Services was obtained through the IBM Quantum Hub at CERN.

### v2

@@ Acknowledgements.
We acknowledge Matteo Robbiati for providing help and support in executing the QNN on the Qibo hardware and Stefano Carrazza for providing us access to the quantum device. We acknowledge financial support from the Spanish Government (Agencia Estatal de Investigación MCIN/AEI/ 10.13039/501100011033) Grant No. PID2020-114473GB-I00, and Generalitat Valenciana Grants No. PROMETEO/2021/071 and ASFAE/2022/009 (Planes Complementarios de I+D+i, NextGenerationEU). This work is also supported by the Ministry of Economic Affairs and Digital Transformation of the Spanish Government and NextGenerationEU through the Quantum Spain project, and by CSIC Interdisciplinary Thematic Platform on Quantum Technologies (PTI-QTEP+). JML is supported by Generalitat Valenciana (Grant No. ACIF/2021/219) and CSIC (Grant No. IMOVE23124). LC is supported by Generalitat Valenciana GenT Excellence Programme (CIDEGENT/2020/011) and ILINK22045. MG and SV are supported by CERN through the CERN QTI. Access to the IBM Quantum Services was obtained through the IBM Quantum Hub at CERN.